a ball rolls off a platform that is 13 meters above the ground. the ball's horizontal velocity as it leaves the platform is 5m/s. using the approximate value of g=10 m/s^2 how much time does it take for the ball to hit the ground?
Answer #1

Solution-

a) Using formula
s = ut + at^2/2
here
s = vertical displacement = -13 m (Here negative indicates final height below initial height)
u = initial vertical velocity = 0 m/s
t = time = To find
a = acceleration by gravity = -9.8 m/s²
Putting the values in the equation
-13 = 0*t + (-10)t^2/2
-13 = -5×t²
t² = 2.6
t = 1.61 s